OpenStudy (keeks):

k+4/4 + k-4/4= k+4/4k @AloneS

OpenStudy (seratul):

Okay, you can add k+4/4 and k-4/4 to get 2k/4=k=4/4k Then cross multiply. Tell me what you get.

OpenStudy (seratul):

2k/4=k+4/4k

alones (alones):

wow x.x im way off..i got no solution LOL

OpenStudy (keeks):

I got it thanks